# Unbundle-Rewrite-Rebundle: Runtime Detection and Rewriting of Privacy-Harming Code in JavaScript Bundles

### v2

@@ Abstract.
This work presents Unbundle-Rewrite-Rebundle (URR), a system for detecting privacy-harming portions of bundled JavaScript code, and rewriting that code at runtime to remove the privacy harming behavior without breaking the surrounding code or overall application. URR is a novel solution to the problem of JavaScript bundles, where websites pre-compile multiple code units into a single file, making it impossible for content filters and ad-blockers to differentiate between desired and unwanted resources. Where traditional content filtering tools rely on URLs, URR analyzes the code at the AST level, and replaces harmful AST sub-trees with privacy-and-functionality maintaining alternatives.
We present an open-sourced implementation of URR as a Firefox extension, and evaluate it against JavaScript bundles generated by the most popular bundling system (Webpack) deployed on the Tranco 10k. We measure the performance, measured by precision (1.00), recall (0.95), and speed (0.43s per-script) when detecting and rewriting three representative privacy harming libraries often included in JavaScript bundles, and find URR to be an effective approach to a large-and-growing blind spot unaddressed by current privacy tools.
@@ 1. Introduction
An enormous body of research has established Web content filtering (e.g., blocking advertising, tracking, and other unwanted network requests on websites) as an important and effective technique for improving privacy(Merzdovnik et al., 2017; Gervais et al., 2017), security(Li et al., 2012; Zarras et al., 2014), and performance(Garimella et al., 2017; Pujol et al., 2015). Most Web content filtering approaches rely on crowd sourced lists of regular-expression-like rules that describe which URLs the browser should load, and which should be blocked.
This approach—broadly, URL based content filtering—works because URLs in practice provide useful and stable information about the resources they map to. In some cases this is because of the text in the URL (e.g., browsers can make a reasonable guess about the purpose of JavaScript returned from a URL like
https://advertising.example/tracker.js), or because experts have manually evaluated the resource returned from a URL and found it to be similarly harmful to users.
However, modern Web development practices make URL based content filtering increasingly difficult. Previously, Web applications were often delivered as a collection of discrete JavaScript files, each fetched independently from their own URL (e.g., /script/library.js, /script/tracker.js ,/script/app.js), which allowed URL-based content filtering tools to easily block some parts of an application, but not others. Increasingly though, developers integrate bundling tools as part of their build and deployment practices, compiling all of the libraries and application code into a single file unit, which is delivered to the browser from a single URL (e.g., /script/bundle.js).
These bundling approaches, inadvertently or otherwise, circumvent URL based content filtering tools. When applications are delivered as a single bundled code unit, URL based filtering tools can no longer block just parts of the application; blocking a site’s JavaScript becomes an all-or-nothing proposition. And since blocking all JavaScript on a page breaks useful functionality on many sites, in practice content filtering tools are reduced to blocking nothing, reintroducing the privacy, security, and performance issues the user wanted to avoid in the first place.
@@
URR is a novel, practical solution to a problem that has been explored by a vein of related Web privacy research. Works like (Amjad et al., 2021) identify that bundled applications are widespread and pose a serious challenge to Web privacy, and (Smith et al., 2022) found that blocking these bundled JavaScript resources often broke the benign, desirable parts of websites. Systems like (Smith et al., 2021) showed that bundled applications could be automatically rewritten to prevent privacy harm, though with expensive precomputation, which rendered practical deployment prohibitive. URR is a first-in-class approach to solving the privacy and security harms caused by bundled JavaScript applications, in a method that is performant and practical.
@@
First, the system must identify known privacy harming code (e.g., the code delivered from /script/tracker.js) within the larger bundled application, in the absence of any information about the URL the code originally came from. This identification must be robust even across the kinds of code modifications and transformations JavaScript bundlers make in their build processes (e.g., minification, dead-code elimination, tree-shaking).
Second, URR must remove known-privacy-harming code from the bundled application, without breaking desireable functionality in the surrounding code. Just deleting unwanted libraries from a bundled application will, in practice, be counter productive, since the application will fail when trying to access now-deleted functions and classes defined by the deleted library. A useful solution must remove the unwanted, target libraries from the bundled application without breaking surrounding code.
Third, such a system must be performant, and be able to Debundle, analyze, modify, and reconstitute bundled Web applications at runtime, and quickly, in a way that maintains the usefulness of the Web application. If the performance overhead of a privacy-and-security preserving system is too costly, then the system is in practice unusable, and so not meaningfully useful at benefit Web users.
URR is implemented in several parts: i. as a database of signatures of ASTs of real world known-privacy-harming code, ii. a library of crowd sourced privacy-preserving alternative implementations of privacy-harming code, designed to remove privacy harming behaviors without impacting surrounding application code, iii. a browser extension that, at runtime, decomposes a bundled JavaScript application into its constituent libraries, detects the sub-ASTs from the bundled application that come from known-privacy-harming libraries, and rewrites the bundled application with the stub-libraries in place of the privacy-harming versinos.
@@
    An open sourced, empirically tuned system for:
@@
      reconstituting the resulting new application into a new bundle, that can then be passed to the browser’s JavaScript engine for normal execution.
@@
    An empirical evaluation of the accuracy and performance of our system when applied to a representative crawl of the Web, finding that our system results in libraries being blocked on 7% of the top 10K sites in the Tranco list, within practical performance bounds.
@@
    An open source implementation of our system as a Firefox extension, along with the complete dataset for all discussed figures and measurements.
@@ 2. Background
This section first provides a primer on concepts relevant to bundling and their use in web development. It then presents a simple example that highlights the limitations of existing content blocking approaches. The section concludes by outlining the properties of an effective solution.
@@ 2.1. Bundles and Relevant Concepts
As websites and web applications grow more complex, they require a plethora of functionality, often aided by numerous libraries and dependencies. Handling these dependencies can prove to be quite strenuous, especially when needing to take into account the range of platforms on which the code needs to correctly execute. JavaScript bundlers are essential tools used by web developers to streamline the handling of code and dependencies within complex web applications. At its core, a JavaScript bundler is a utility that gathers and wraps code from multiple JavaScript files. Bundles not only reduce the number of network requests required to load a web page but also optimize the handling of dependencies and various aspects of development and production environments. The simplicity of the example hides certain nuances necessary to understand the complexity of bundles. Below, we introduce a few fundamental concepts that can help better understand JavaScript bundles.
@@
Web developers design and create websites in different environments than browsers. These environments have their own caveats and follow different programming concepts and philosophies. Modular programming is a general programming concept where developers separate complex application code into independent pieces called modules. A module forms the atomic unit of a bundle, and rougly corresponds to a code snippet relevant to a file or library that exports functionality that is consumed by other modules. Popular JavaScript development environments like Node.js (Nod, 2009) adopt a modular programming approach. Developers can create their own modules and use reuse modules available in the npm registry (NPM, 2010).
@@
Like development environments and programming approaches, JavaScript module systems are also not a monolith. Depending on the context in which they are consumed and executed, JavaScript modules express functionality in different ways. The two most popular module systems for JavaScript are (1) the ECMAScript modules (ESM) which are consumed with import statements, and (2) CommonJS (CJS) modules which are consumed with require statements. While Node.js supports both types of modules, web browsers only recognize import statements. Bundles therefore include wrappers around modules and provide workarounds for require statements within web browsers.
@@
Larger projects include multiple libraries and packages, and as a result comprise numerous, inter-dependent modules. When bundles gather and parse all the modules that need to be combined, they create a dependency graph that helps determine (1) the order and chain of dependencies between modules, and (2) which code snippets can be combined within a module and which snippets need to be split across multiple modules.
@@
During development, bundles provide source maps as a key to reverse minification and debug parts of code. These files help map minified code back to their unminified counterparts, and hence identify individual modules. However, source maps are not available by default in production, making it difficult to reverse engineer bundled code in the wild.
@@
Examples of popular JavaScript bundlers include Webpack (Web, 2014c), Browserify (Bro, 2020), Rollup (Rol, 2015), and Parcel (Par, 2018), each of which offer unique features and advantages. In this work, we focus on Webpack because it is the most popular and mature JavaScript bundler as determined from GitHub stars (Web, 2014a), NPM weekly downloads (Web, 2014b), and prior work (Rack and Staicu, 2023a).
@@ 2.2. Motivating Example
In this section we present an example that shows how typical content blocking approaches work and why they are ineffective against bundles.
@@
Consider the toy example presented on the left half of Figure 1. Here, the websites loads two scripts, each of which define global functions. The first script, setup.js, loads from the website’s domain itself and contains benign code relevant to the website’s functionality. The second script, track.js, loads from a known tracking domain, tracker.com, and returns a script that creates a unique identifier to track the user. When the website includes the two scripts from two separate <script> tags, it triggers two network requests, one to website.com and another to tracker.com.
@@
A typical approach to blocking privacy-harming scripts involves the use of a curated list of domains and regular expressions (for example, from EasyList (EasyList Authors, 2023), EasyPrivacy (EasyPrivacy Authors, 2023)). These lists are developed from manual contributions and include domains and paths to known privacy-harming resources. Content blocking tools (e.g., AdBlock (eyeo GmbH, 2023), uBlock (Hill and Rolls, 2023)) pull from these filter lists. The tools intercept outgoing network requests, compare them against entries in filter lists, and create interventions if they find a match. In the toy example, a filter list contributor might add the rule ||tracker.com/track.js. Thereafter, when the website creates two network requests, the content-blocking tool permits a request to script.js but blocks the request to tracker.com/track.js. This way, the privacy-harming script is neither loaded nor executed in the user’s browsing session.
@@
Consider the scenario presented in the right half of Figure 1. Unlike the previous example, the website includes a single <script> tag that fetches code from its own server. The resulting network request does not have a corresponding entry in the filter list and is therefore not blocked by the content blocking tool. The server responds with a bundled script that includes both, benign code (setup()) and privacy-harming code (track()). Bundled resources expose multiple shortcomings of existing content-blocking approaches, which we briefly describe below.
  • Privacy-harming code can be fetched from multiple, variable resource paths, including from first-party domains, making it impossible for contributors to manual detect and curate a comprehensive, non-exhaustive list of entries.
  • Bundles include code from multiple resources. As a result, privacy-harming code is embedded along with necessary and benign functionality. Blocking the network request itself can break the website. Existing approaches need to be adopt an approach that targets embedded resources instead of the entire resource itself.
@@ 2.3. Properties of an Ideal Solution
First, a robust solution should be able to identify and target specific modules. Unlike typical content blocking approaches, the solution cannot rely on domain-based blocking or attempt to replace the a script in its entirety.
Second, as a corollary to the previous point, the solution should have limited impact on benign functionality, i.e., the solution should limit the effect of its intervention to privacy-harming code, leaving execution of other modules untouched. This also involves ensuring that any dependancies on the targeted module are handled in a way that limits side-effects.
Third, the solution should be generalizable across multiple dimensions. It should be applicable to multiple privacy-harming libraries (and multiple versions of libraries) that are of interest to content blocking tools. Additionally, it should also be generalizable to different bundling strategies and robust against minification and obfuscation techniques.
@@ 3. Unbundle-Rewrite-Rebundle Design
Unbundle-Rewrite-Rebundle (URR) adopts a static analysis approach that leverages the code structures of privacy-harming modules to identify and neutralize them when embedded in bundled scripts without disrupting the functionality of other components of the application.
URR adopts a four-step process (see Figure 2):
  • First, URR generates an Abstract Syntax Tree (AST) representation of a given script. It then analyzes the structure to identify if the script is a bundle that comprises one or more modules. If so, URR gathers the component modules for further anlaysis.
@@
  • Finally, URR replaces each marked privacy-harming module with a corresponding, benign replacement. In doing so, URR ensures that access and use of the replacement does not break other parts of website functionality. It then stitches the bundled script back together and moves it along for consumption and execution.
All the steps mentioned above are performed in real-time, i.e., when resources are loaded in the user’s browser. However, processed representations of target modules and benign replacements are created and gathered offline.
@@ 3.1. Gathering Modules
Thereafter, URR uses the AST representation to determine two aspects of the loaded script. First, URR determines if the script comprises a bundle that includes one or more modules. Second, URR gathers the sub-trees corresponding to each of the identified modules.
Below, we describe the webpack-specific implementation for this phase.

### v3

@@ Abstract.
This work presents Unbundle-Rewrite-Rebundle (URR), a system for detecting privacy-harming portions of bundled JavaScript code and rewriting that code at runtime to remove the privacy-harming behavior without breaking the surrounding code or overall application. URR is a novel solution to the problem of JavaScript bundles, where websites pre-compile multiple code units into a single file, making it impossible for content filters and ad-blockers to differentiate between desired and unwanted resources. Where traditional content filtering tools rely on URLs, URR analyzes the code at the AST level, and replaces harmful AST sub-trees with privacy-and-functionality maintaining alternatives.
We present an open-sourced implementation of URR as a Firefox extension and evaluate it against JavaScript bundles generated by the most popular bundling system (Webpack) deployed on the Tranco 10k. We evaluate URR by precision (1.00), recall (0.95), and speed (0.43s per script) when detecting and rewriting three representative privacy-harming libraries often included in JavaScript bundles, and find URR to be an effective approach to a large-and-growing blind spot unaddressed by current privacy tools.
Web privacy; Content blocking
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690262isbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Web application security
@@ 1. Introduction
An enormous body of research has established Web content filtering (e.g., blocking advertising, tracking, and other unwanted network requests on websites) as an important and effective technique for improving privacy(Merzdovnik et al., 2017; Gervais et al., 2017), security(Li et al., 2012; Zarras et al., 2014), and performance(Garimella et al., 2017; Pujol et al., 2015). Most Web content filtering approaches rely on crowd-sourced lists of regular-expression-like rules that describe which URLs the browser should load and which should be blocked.
This approach—broadly, URL-based content filtering—works because URLs, in practice, provide useful and stable information about the resources they map to. In some cases, this is because of the text in the URL (e.g., browsers can make a reasonable guess about the purpose of JavaScript returned from a URL like (https://advertising
.example/tracker.js
), or because experts have manually evaluated the resource returned from a URL and found it to be similarly harmful to users.
However, modern Web development practices make URL-based content filtering increasingly difficult. Previously, Web applications were often delivered as a collection of discrete JavaScript files, each fetched independently from their own URL (e.g., /script/library.js, /script/tracker.js,/script/app.js), which allowed URL-based content filtering tools to easily block some parts of an application, but not others. Increasingly, though, developers integrate bundling tools as part of their build and deployment practices, compiling all of the libraries and application code into a single file unit, which is delivered to the browser from a single URL (e.g., /script/bundle.js).
These bundling approaches, inadvertently or otherwise, circumvent URL-based content filtering tools. When applications are delivered as a single bundled code unit, URL-based filtering tools can no longer block just parts of the application; blocking a site’s JavaScript becomes an all-or-nothing proposition. And since blocking all JavaScript on a page breaks useful functionality on many sites, in practice, content filtering tools are reduced to blocking nothing, reintroducing the privacy, security, and performance issues the user wanted to avoid in the first place.
@@
URR is a novel, practical solution to a problem that a vein of related Web privacy research has explored. Works like (Amjad et al., 2021) identify that bundled applications are widespread and pose a serious challenge to Web privacy, and (Smith et al., 2022) found that blocking these bundled JavaScript resources often broke the benign, desirable parts of websites. Systems like (Smith et al., 2021) showed that bundled applications could be automatically rewritten to prevent privacy harm, though with expensive precomputation, which rendered practical deployment prohibitive. URR is a first-in-class approach to solving the privacy and security harms caused by bundled JavaScript applications in a performant and practical method.
@@
First, the system must identify known privacy-harming code (e.g., the code delivered from /script/tracker.js) within the larger bundled application, without any information about the URL from which the code originally came. This identification must be robust even across the kinds of code modifications and transformations JavaScript bundlers make in their build processes (e.g., minification, dead-code elimination, tree-shaking).
Second, URR must remove known-privacy-harming code from the bundled application without breaking desirable functionality in the surrounding code. Just deleting unwanted libraries from a bundled application will, in practice, be counter-productive since the application will fail when trying to access now-deleted functions and classes defined by the deleted library. A useful solution must remove the unwanted, target libraries from the bundled application without breaking surrounding code.
Third, such a system must be performant and be able to Unbundle, analyze, modify, and reconstitute bundled Web applications at runtime, and quickly, in a way that maintains the usefulness of the Web application. If the performance overhead of a privacy-and-security preserving system is too costly, then the system is, in practice, unusable and so not meaningfully useful to benefit users.
URR is implemented in several parts: i. as a database of signatures of ASTs of real-world known-privacy-harming code, ii. a library of crowd-sourced privacy-preserving alternative implementations of privacy-harming code designed to remove privacy harming behaviors without impacting the surrounding application code, iii. a browser extension that, at runtime, decomposes a bundled JavaScript application into its constituent libraries, detects the sub-ASTs from the bundled application that come from known-privacy-harming libraries and rewrite the bundled application with the stub libraries in place of the privacy-harming versions.
@@
    An open-sourced, empirically tuned system for:
@@
      reconstituting the resulting new application into a new bundle that can then be passed to the browser’s JavaScript engine for normal execution.
@@
    An empirical evaluation of the accuracy and performance of our system when applied to a representative crawl of the Web, finding that our system results in libraries being blocked on 7% of the top 10K sites in the Tranco list within practical performance bounds.
@@
    An open-source implementation of our system as a Firefox extension, along with the complete dataset for all discussed figures and measurements, available at https://github.com/masood/urr.
@@ 2. Background
This section first provides a primer on concepts relevant to bundling and their use in web development. It then presents a simple example that highlights the limitations of existing content-blocking approaches. The section concludes by outlining the properties of an effective solution.
@@ 2.1. Bundles and Relevant Concepts
As websites and web applications grow more complex, they require a plethora of functionality, often aided by numerous libraries and dependencies. Handling these dependencies can prove to be quite strenuous, especially when considering the range of platforms on which the code needs to execute correctly. JavaScript bundlers are essential tools used by web developers to streamline the handling of code and dependencies within complex web applications. At its core, a JavaScript bundler is a utility that gathers and wraps code from multiple JavaScript files. Bundles not only reduce the number of network requests required to load a web page but also optimize the handling of dependencies and various aspects of development and production environments. The simplicity of the example hides certain nuances necessary to understand the complexity of bundles. Below, we introduce a few fundamental concepts that can help better understand JavaScript bundles.
@@
Web developers design and create websites in different environments than browsers. These environments have their own caveats and follow different programming concepts and philosophies. Modular programming is a general programming concept where developers separate complex application code into independent pieces called modules. A module forms the atomic unit of a bundle and roughly corresponds to a code snippet relevant to a file or library that exports functionality that is consumed by other modules. Popular JavaScript development environments like Node.js [^1] https://nodejs.org/en/ adopt a modular programming approach. Developers can create their own modules and use reuse modules available in the npm registry [^2] https://www.npmjs.com/.
@@
Like development environments and programming approaches, JavaScript module systems are also not a monolith. Depending on the context in which they are consumed and executed, JavaScript modules express functionality in different ways. The two most popular module systems for JavaScript are (1) the ECMAScript modules (ESM), which are consumed with import statements, and (2) CommonJS (CJS) modules, which are consumed with require statements. While Node.js supports both types of modules, web browsers only recognize import statements. Bundles, therefore, include wrappers around modules and provide workarounds for require statements within web browsers.
@@
Larger projects include multiple libraries and packages and, as a result, comprise numerous interdependent modules. When bundles gather and parse all the modules that need to be combined, they create a dependency graph that helps determine (1) the order and chain of dependencies between modules and (2) which code snippets can be combined within a module and which snippets need to be split across multiple modules.
@@
During development, bundles provide source maps as a key to reverse minification and debug parts of code. These files help map minified code back to their unminified counterparts and, hence, identify individual modules. However, source maps are unavailable by default in production, making it difficult to reverse engineer bundled code in the wild.
@@
Examples of popular JavaScript bundlers include Webpack [^3] https://webpack.js.org/, Browserify [^4] https://browserify.org/, Rollup [^5] https://rollupjs.org/, and Parcel  [^6] https://parceljs.org/, each of which offer unique features and advantages. In this work, we focus on Webpack because it is the most popular and mature JavaScript bundler as determined from GitHub stars [^7] https://github.com/webpack/webpack, NPM weekly downloads [^8] https://www.npmjs.com/package/webpack, and prior work (Rack and Staicu, 2023).
@@ 2.2. Motivating Example
In this section, we present an example that shows how typical content-blocking approaches work and why they are ineffective against bundles.
@@
Consider the toy example presented on the left half of Figure 1. Here, the website loads two scripts, each of which defines global functions. The first script, setup.js, loads from the website’s domain itself and contains benign code relevant to the website’s functionality. The second script, track.js, loads from a known tracking domain, tracker.com, and returns a script that creates a unique identifier to track the user. When the website includes the two scripts from two separate <script> tags, it triggers two network requests, one to website.com and another to tracker.com.
@@
A typical approach to blocking privacy-harming scripts involves the use of a curated list of domains and regular expressions (for example, from EasyList, EasyPrivacy [^9] https://easylist.to/). These lists are developed from manual contributions and include domains and paths to known privacy-harming resources. Content blocking tools (e.g., AdBlock, uBlock) pull from these filter lists. The tools intercept outgoing network requests, compare them against entries in filter lists, and create interventions if they find a match. In the toy example, a filter list contributor might add the rule ||tracker.com/track.js. Thereafter, when the website creates two network requests, the content-blocking tool permits a request to script.js but blocks the request to tracker.com/track.js. This way, the privacy-harming script is neither loaded nor executed in the user’s browsing session.
@@
Consider the scenario presented in the right half of Figure 1. Unlike the previous example, the website includes a single <script> tag that fetches code from its own server. The resulting network request does not have a corresponding entry in the filter list and is, therefore, not blocked by the content-blocking tool. The server responds with a bundled script that includes both benign code (setup()) and privacy-harming code (track()). Bundled resources expose multiple shortcomings of existing content-blocking approaches, which we briefly describe below.
  • Privacy-harming code can be fetched from multiple, variable resource paths, including from first-party domains, making it impossible for contributors to manually detect and curate a comprehensive, non-exhaustive list of entries.
  • Bundles include code from multiple resources. As a result, privacy-harming code is embedded along with necessary and benign functionality. Blocking the network request itself can break the website. Existing approaches must adopt an approach that targets embedded resources instead of the entire resource itself.
@@ 2.3. Properties of an Ideal Solution
First, a robust solution should be able to identify and target specific modules. Unlike typical content-blocking approaches, the solution cannot rely on domain-based blocking or attempt to replace the script entirely.
Second, as a corollary to the previous point, the solution should have limited impact on benign functionality, i.e., the solution should limit the effect of its intervention to privacy-harming code, leaving execution of other modules untouched. This also involves ensuring that any dependencies on the targeted module are handled in a way that limits side effects.
Third, the solution should be generalizable across multiple dimensions. It should apply to multiple privacy-harming libraries (and multiple versions of libraries) that interest content-blocking tools. Additionally, it should also be generalizable to different bundling strategies and robust against minification and obfuscation techniques.
@@ 3. Unbundle-Rewrite-Rebundle Design
Unbundle-Rewrite-Rebundle (URR) adopts a static analysis approach that leverages the code structures of privacy-harming modules to identify and neutralize them when embedded in bundled scripts without disrupting the functionality of other application components. URR adopts a four-step process (see Figure 2):
  • First, URR generates an Abstract Syntax Tree (AST) representation of a given script. It then analyzes the structure to identify if the script is a bundle that comprises one or more modules. If so, URR gathers the component modules for further analysis.
@@
  • Finally, URR replaces each marked privacy-harming module with a corresponding benign replacement. In doing so, URR ensures that access and use of the replacement does not break other parts of website functionality. It then stitches the bundled script and moves it along for consumption and execution.
All the steps mentioned above are performed in real-time, i.e. when resources are loaded in the user’s browser. However, processed representations of target modules and benign replacements are created and gathered offline.
@@ 3.1. Gathering Modules
Thereafter, URR uses the AST representation to determine two aspects of the loaded script. First, URR determines if the script comprises a bundle with one or more modules. Second, URR gathers the sub-trees corresponding to each identified module.
Below, we describe the webpack-specific implementation.